# LocMoE: A Low-overhead MoE for Large Language Model Training

### v1

@@ 3.1 PanGu-
The PanGu- architecture consists of both dense and sparse Transformer encoder layers, stacked Transformer decoder layers modeled in the autoregressive language, and a query layer. The general knowledge is learned by dense layers on the bottom, while domain-specific knowledge is acquired by the expert network. The sparse Transformer layer of PanGu-, with several conditionally activated feedforward sublayers, incorporates the MoE principle, as displayed in Figure 3. The RRE module is responsible for routing the token to the appropriate expert. It contains two levels of routing: in the first level, the experts are grouped by domains, and the token is assigned to one of the groups. In the second level, the token is routed to a particular expert of this group homogeneously. The second level of routing can be viewed as random hash routing, which does not contain learnable parameters.
@@ 3.2 MoE with Local Routing Strategy
The introduction of expert capacity aims to avoid the training block caused by the assignment imbalance of tokens. In general, an empirical expert capacity factor is set to limit the scale of the expert capacity: . The computational workload of experts can be equalized in this way. However, the work of pMoE reveals that the sample size each expert needs to process has its lower bound [?], which can provably reduce the training costs. Inspired by the work on pMoE, we migrated the assumptions of data distribution in MoE from CV to the NLP domain, in conjunction with the network structure, while further discovering some significant conclusions:
@@
Suppose and tokens with size of uniformly distributed on unit sphere with dimensions.
@@
On account of the spherical symmetry, the probabilities for are equivalent for all experts under the conditions of orthogonal gating weights (brought from the GAP layer). That is, .
@@ 4.1 Analysis for expert capacity
To verify our proof on Lemma 1, the angle between tokens, as well as the angle between the token and the gating weight, are explored, shown in Figure 8 and 9, respectively. Figure 8 at row i and conlumn j plots the distribution of cosine similarities between every pair of tokens routed to the expert and , respectively. Specifically, the diagonal ones denote the distributions of tokens from the same expert. It can be seen that the tokens routed to the same expert are more alike, with the cosine similarity closer to 1.
@@ 4.2 Ablation Analysis
The aim of this paper is to explore the technique that can improve the training efficiency of MoE. Therefore, the ablation study is built around aspects including the proportion of computation and communication time, load equalization, and astringency. Moreover, in order to prove that the modifications do not affect the model accuracy, the results for inference are also evaluated for verification.
@@
Taking the cluster of 64N as an instance, Figure 12 portrays the distribution of expert assignments in different MoEs during the training process. The token distributions are plotted every 100 epochs, and the subfigures for the first 200 epochs are omitted. Since HashMoE adopts an absolute balance strategy, the allocation of tokens is quite balanced at initialization. However, SwitchMoE and LocMoE initialize from the allocation to a single expert. To avoid misinterpretation, the analysis begins at epoch 200. The vertical axis indicates the number of tokens assigned to each communication group, and the horizontal axis indicates the index of experts. There are 16 experts in our experiments, and the index range is from 0 to 15. The cumulative number of tokens routed to expert can be observed along the specific vertical axis corresponding to the expert. Each occurrence of a non-zero value means a new token being assigned to this expert. Successive color bars indicate that shuffled tokens are continuously assigned to the same experts, thus causing imbalances to arise. As can be seen from the figure, the rigid constraints in HashMoE ensure that its assignment is even. However, almost no token is routed to experts with an index of 9 to 15 in the subfigure of SwitchMoE. It results in nearly 40% of the experts’ invalidation; to make matters worse, the phenomena of ”winner-take-all” is pronounced in expert number 5 and expert number 6. LocMoE, due to the localized bootstrapping, can allocate the token to these experts evenly during the training process, indicating that the dual constraints of auxiliary and locality loss can steadily enhance resource utilization.
@@ A.3 Proof for Theorem 1
Refer to the assumption about distributions of class-discriminative and class-irrelevant patterns in pMoE [?], with analogy, the tokens satisfy can be regarded as the class-discriminative token. Then, the problem we need to explore can be converted to find the minimum amount of tokens that make at least one class-discriminative token routed to expert . Suppose is the probability that the token routed to the expert is a class-discriminative token; we have:
@@ C. Experimental Environment
Hyperparameter | Description | Value |
adam_eps |
Terms to increase the stability of numerical calculations |
1e-6 |
batch_size |
The size of data input to the model for training each time, related to the number of devices |
32 |
expert_num_per_dp_dim |
Number of experts per communication group |
1 |
expert_parallel |
Number of experts in parallel |
16 |
moe_layer_num |
Number of MoE layers |
8 |
num_heads |
Number of parallel heads |
40 |
op_level_model_parallel_num |
Number of parallel models |
8 |
sink_size |
The size of data executed per sink |
16 |
@@ E. Measurement Metric
The examples of NLP task samples are displayed in Figure 16 and Figure 17. These two tasks are both from the Huawei Certified ICT Expert (HCIE) test. It can be seen that LocMoE is better at learning and understanding domain knowledge and analyzing the answers in a more specific way.

### v2

@@ 3.1 PanGu-
The PanGu- architecture consists of both dense and sparse Transformer encoder layers, stacked Transformer decoder layers modeled in the autoregressive language, and a query layer. The sparse Transformer layer of PanGu-, with several conditionally activated feedforward sublayers, incorporates the MoE principle, as displayed in Figure 3. The RRE module is responsible for routing the token to the appropriate expert. It contains two levels of routing: in the first level, the experts are grouped by domains, and the token is assigned to one of the groups. In the second level, the token is routed to a particular expert of this group homogeneously. The second level of routing can be viewed as random hash routing, which does not contain learnable parameters.
@@ 3.2 MoE With Local Routing Strategy
The introduction of expert capacity aims to avoid the training block caused by the assignment imbalance of tokens. In general, an empirical expert capacity factor is set to limit the scale of the expert capacity: . , , , and denote the batch size, the capacity factor, the degree of expert parallelism, and the number of experts, respectively. The computational workload of experts can be equalized in this way. However, the work of pMoE reveals that the sample size each expert needs to process has its lower bound [?], which can provably reduce the training costs. Inspired by the work on pMoE, we migrated the assumptions of data distribution in MoE from CV to the NLP domain, in conjunction with the network structure, while further discovering some significant conclusions:
@@
Suppose all tokens with size of uniformly distributed on the unit sphere, that is, .
@@
Suppose denotes the expert to which the token is routed. On account of the spherical symmetry, the probabilities for are equivalent for all experts under the conditions of orthogonal gating weights (brought from the GAP layer). That is, .
@@ 4.1 Analysis for Expert Capacity
To verify our proof on Lemma 1, the angle between tokens, as well as the angle between the token and the gating weight, are explored, shown in Figure 8 and 9, respectively. Figure 8 at row and conlumn plots the distribution of cosine similarities between every pair of tokens routed to the expert and , respectively. Specifically, the diagonal ones denote the distributions of tokens from the same expert. It can be seen that the tokens routed to the same expert are more alike, with the cosine similarity closer to 1.
@@ 4.2 Ablation Analysis
The ablation study is built around aspects including the proportion of computation and communication time, load equalization, and astringency. Moreover, in order to prove that the modifications do not affect the model accuracy, the results for inference are also evaluated for verification.
@@
Taking the cluster of 64N as an instance, Figure 12 portrays the distribution of expert assignments in different MoEs during the training process. Since HashMoE adopts an absolute balance strategy, the allocation of tokens is quite balanced at initialization. However, SwitchMoE and LocMoE initialize from the allocation to a single expert. To avoid misinterpretation, the analysis begins at epoch 200. The vertical axis indicates the number of tokens assigned to each communication group, and the horizontal axis indicates the index of experts. There are 16 experts in our experiments, and the index range is from 0 to 15. The cumulative number of tokens routed to expert can be observed along the specific vertical axis corresponding to the expert. Each occurrence of a non-zero value means a new token being assigned to this expert. Successive color bars indicate that shuffled tokens are continuously assigned to the same experts, thus causing imbalances to arise. As can be seen from the figure, the rigid constraints in HashMoE ensure that its assignment is even. However, almost no token is routed to experts with an index of 9 to 15 in the subfigure of SwitchMoE. It results in nearly 40% of the experts’ invalidation; to make matters worse, the phenomena of ”winner-take-all” is pronounced in expert number 5 and expert number 6. LocMoE, due to the localized bootstrapping, can allocate the token to these experts evenly during the training process, indicating that the dual constraints of auxiliary and locality loss can steadily enhance resource utilization.
@@ A.2 Proof for Lemma 3
Hyperparameter | Description | Value |
adam_eps | Terms to increase the stability of numerical calculations | 1e-6 |
batch_size | The size of data input to the model for training each time, related to the number of devices | 32 |
expert_num_per_dp_dim | Number of experts per communication group | 1 |
expert_parallel | Number of experts in parallel | 16 |
moe_layer_num | Number of MoE layers | 8 |
num_heads | Number of parallel heads | 40 |
op_level_model_parallel_num | Number of parallel models | 8 |
sink_size | The size of data executed per sink | 16 |
@@ A.3 Proof for Theorem 1
Refer to the assumption about distributions of class-discriminative and class-irrelevant patterns in pMoE [?], with analogy, the tokens satisfy can be regarded as the class-discriminative token. Then, the problem we need to explore can be converted to find the minimum amount of tokens that make at least one class-discriminative token routed to expert .
Suppose is the probability that the token routed to the expert is a class-discriminative token; we have:
@@ E. Measurement Metric
Contribution Statement
Jing Li, Zhijie Sun, and Xuan He wrote this paper and contributed equally to this work. Li Zeng, Yi Lin, Entong Li, and Binfan Zheng provided experimental analysis and offered suggestions for this paper. Rongqian Zhao and Xin Chen are the project leaders and provide support for this work.
Acknowledgments
We thank all anonymous reviewers for their valuable feedback. We would like to express our appreciation to Dachao Lin (Dr. Lin) for the improvements in theoretical proof in this paper. We are grateful for the guidance from the Noah’s Ark Lab on model training. Moreover, the contributions of the MindSpore team and employees participating in model evaluation are greatly appreciated.